Consider a manufacturing line with 3 robotic machines used to create a widget. Each robotic machine either performs (P) perfectly or creates a (D) defect. Denote the outcome of the widget by creating 3-tuples of P's and D's.

1) Find the sample space S.

2) Find the event A that at least one robotic machine created a defect (D).

3) Find event B that all three robotic machines performed the same way.

4) Are the events A and B mutually exclusive? Show mathematically using set theory notation.
Answer #1

A robotic machine performs perfectly is denoted by P

A robotic machine creates a defect is given by D

There are 3 robotic machines

The sample space will contain 23 =8 elements

The sample space is given by

s={PPP,PPD,PDP,PDD,DPP,DPD,DDP,DDD}

A be the event that at least one robotic machine created a defect

A={PPD,PDP,PDD,DPP,DPD,DDP,DDD}

n(A)=no. of cases favourable to the event=7

n=no. of possible outcomes in the sample space s=8

so P(A)= probability that at least one robotic machine created a defect=n(A)/n=7/8

now B be the even t that all three robotic machines performed the same way

B={PPP,DDD}

n(B)=no. of cases favourable to the event=2

P(B)=probability that all the machines performed in the same way=n(B)/n=2/8=.25

now we know that two events are mutually exclusive if

P(A\capB)=P(A)*P(B)

Now A\capB= {DDD}

n(A\capB)=no. of cases favourable to the event A\capB=1

P(A\capB)=n(A\capB)/n=1/8

P(A)*P(B)=7/8*2/8=14/64

so P(A\capB)\neqP(A)*P(B)

so the events A and B are not mutually exclusive.

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image